OpenStudy (anonymous):

6.246/X x 100 = 5% it's 6.246 over X times 100 = 5 %, what is X?

OpenStudy (anonymous):

yes, it's like in the pic

OpenStudy (phi):

You can use the equation editor (see button on the lower left below the input area. is this what you mean \[ \frac{6.246}{x}⋅100=5\% \]

OpenStudy (phi):

change 5% to 0.05 or 5/100 multiply both sides by x you get \[ 6.246 \cdot 100 = 0.05 x \] simplify the left side \[ 624.6 = 0.05 x\] divide both sides by 0.05

OpenStudy (anonymous):

why would you change that to .005 if the times 100 was already there to change it into a percent?

OpenStudy (phi):

I was solving what you wrote. Is the question is 6.246 is 5% of what number ?

OpenStudy (anonymous):

no. i was just given the equation written, no question.

OpenStudy (anonymous):

I got 124.92 as answer

OpenStudy (phi):

if 124.92 is the answer, then it is clearer to say \[ \frac{6.246}{x}= 0.05 \]

OpenStudy (anonymous):

yeah I think it's the same as that

OpenStudy (phi):

you could write it \[ \frac{6.246}{x} = 5\% \] but that is not as clear you could also write it as \[ \frac{6.246}{x}\cdot 100 = 5 \]